OpenStudy (anonymous):

22/3x+1/9y=2 whats thge slope

OpenStudy (anonymous):

slope=-66

OpenStudy (anonymous):

u meant the equation: 22x/3+y/9=2 rite?

OpenStudy (anonymous):

if its the case where, x and y are in denominators, then u differentiate wrt x and re-arrange to get dy/dx,which is the slope.

OpenStudy (anonymous):

noooo,,2/3x+1/9y=2

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

differentiate the equation wrt x: we get -2/3x^2-(1/9y^2)dy/dx=0 this gives us, dy/dx=-2*9y^2/3x^2=-6y^2/x^2